OpenStudy (anonymous):

@Beginning Algebra find 2 consecutive odd integers whose sum is 160. @Beginning Algebra

OpenStudy (anonymous):

Consecutive odds are the same as consecutive evens... 1,3,5,7,9... you add 2 to get to the next number so the first two consecutive odds are x and x + 2 so x + x + 2 = 160 2x + 2 = 160 2x = 158 x = 79 so the numbers are 79 and 81

OpenStudy (asnaseer):

lets call the first odd integer x. Then the next one along must be x+2. So \[x + (x+2) = 160\] \[x + x + 2 = 160\] \[2x+2 = 160\] \[2x = 158\] \[x = 79\] So the two consecutive inteers must be 79 and 81

OpenStudy (agreene):

I'd like to point out the odd counting function: \[N_{odd}=2k-1\]